### v2

@@ 3 Numerical Test
For the convergence test we set $\Omega=(0,1)^{2}$ which is identified with the two-torus $\mathbb{T}^{2}$. This accounts for the periodic boundary conditions. We set $g^{*}=g^{n}$ and solve the nonlinear system with a Newton method with tolerance $10^{-12}$ in the $L^{2}$-norm. We consider the initial data
$\phi_{(}x,y)$ $=0.4+0.2\sin(2\pi x)^{2}\sin(2\pi y),\qquad\theta_{0}(x,y)=1.+0.2\sin(2\pi x)\sin(2\pi y)$
@@
$e$ $=\tfrac{1}{\theta}+2\phi^{2}(1-\phi)^{2},\qquad s=1-\log(\theta)+1\phi^{2}(1-\phi)^{2}-\tfrac{\gamma}{2}|\nabla\phi|^{2},$
@@
Since no analytical solution is available, the discretization error is estimated by comparing the computed solutions $(\phi_{h,\tau},\mu_{\rho,h,\tau},\theta_{h,\tau},\mathbf{u}_{h,\tau},p_{h,\tau})$ with those computed on uniformly refined grids, here halving space and time discretisation parameter i.e. $(\phi_{h/2,\tau/2},\mu_{\rho,h/2,\tau/2},\theta_{h/2,\tau/2},\mathbf{u}_{h/2,\tau/2},p_{h/2,\tau/2})$. The error quantities for the fully-discrete scheme are given in the energy-norm, i.e.
@@
We observe in Table 1 at least first order convergence, second order for squared norms, in all norms, expect for the gradients of the velocity. Indeed, due to the piecewise quadratic elements and the Crank-Nicolson type approximation of the velocity, this can be expected if the coupling terms do not pollute the convergence rate too much. In general we would however expect that only first order convergence is valid.
@@ 4 Conclusion
In this work we have derived a fully discrete finite element scheme for the non-isothermal Cahn-Hilliard-Navier-Stokes system (1) – (3). To derive the scheme we first formulate the continuous equations into a suitable variational formulation, which allowed for a streamlined Galerkin approximation in space. The time discretisation utilises the usual convex-concave splitting for the Cahn-Hilliard component and a nonlinear implicit time discretisation for the internal energy equation. In future work, we will benchmark the scheme and consider the error analysis. Future extensions to a phase-field dependent density in the spirit of AGG will also be considered.
@@ Acknowledgements.
Support by the German Science Foundation (DFG) via SPP 2256: Variational Methods for Predicting Complex Phenomena in Engineering Structures and Materials (project BR 7093/1-2) and via TRR 146: Multiscale Simulation Methods for Soft Matter Systems (project C3) is gratefully acknowledged. The authors would like to thank D. Trautwein for careful proof-reading of the manuscript.

### v3

@@ 3 Numerical Test
For the convergence test, we set $\Omega=(0,1)^{2}$ which is identified with the two-torus $\mathbb{T}^{2}$. This accounts for the periodic boundary conditions. We set $g^{*}=g^{n}$ and solve the nonlinear system by the Newton method with tolerance $10^{-12}$ in the $L^{2}$-norm. We consider the initial data
$\phi_{(}x,y)$ $=0.4+0.2\sin(2\pi x)\sin(2\pi y),\qquad\theta_{0}(x,y)=1.+0.2\sin(2\pi x)\sin(2\pi y)$
@@
$e$ $=\tfrac{1}{\theta}+2\phi^{2}(1-\phi)^{2},\qquad s=1-\log(\theta)+\phi^{2}(1-\phi)^{2}-\tfrac{\gamma}{2}|\nabla\phi|^{2},$
@@
Since no analytical solution is available, the discretisation error is estimated by comparing the computed solutions $(\phi_{h,\tau},\mu_{h,\tau},\theta_{h,\tau},\mathbf{u}_{h,\tau},p_{h,\tau})$ with those computed on uniformly refined grids, here halving space and time discretisation parameter i.e. $(\phi_{h/2,\tau/2},\mu_{h/2,\tau/2},\theta_{h/2,\tau/2},\mathbf{u}_{h/2,\tau/2},p_{h/2,\tau/2})$. The error quantities for the fully discrete scheme are given in the energy norm, i.e.
@@
We observe in Table 1 at least first order convergence, second order for squared norms, in all norms, except for the gradients of the velocity. Indeed, due to the piecewise quadratic elements and the Crank-Nicolson type approximation of the velocity, this can be expected if the coupling terms do not pollute the convergence rate too much. In general, we would however expect that only first order convergence is valid.
@@ 4 Conclusion
In this work, we have derived the fully discrete finite element scheme for the non-isothermal Cahn-Hilliard-Navier-Stokes system (1) – (3). To derive the scheme we have first formulated the continuous equations in a suitable variational formulation, which allowed for a streamlined Galerkin approximation in space. The time discretisation utilises the usual convex-concave splitting for the Cahn-Hilliard component and the nonlinear implicit time discretisation for the internal energy equation. In future work, we will benchmark the scheme and consider the error analysis. Future extensions to a phase-field dependent density in the spirit of AGG will also be considered.
@@ Acknowledgements.
Support by the German Science Foundation (DFG) via SPP 2256: Variational Methods for Predicting Complex Phenomena in Engineering Structures and Materials (project BR 7093/1-2) and via TRR 146: Multiscale Simulation Methods for Soft Matter Systems (project C3) is gratefully acknowledged. The authors would like to thank D. Trautwein for carefully proofreading the manuscript. We thank the anonymous reviewers for their insightful comments and suggestions.